OpenStudy (anonymous):

The hypotenuse of a right triangle is 17. The lengths of the two legs are whole numbers. What are they? please help thank you

OpenStudy (wolf1728):

I figured out just by using trial and error. By the Pythagorean Theorem, we know that the hypotenuse squared must equal the sum of the squares of the other 2 sides. 17^2 = 289 So, side 1 squared + side 2 squared must equal 289

OpenStudy (anonymous):

ok thank you

OpenStudy (wolf1728):

I figured out the 2 sides just using trial and error they are 8 and 15 8^2 = 64 and 15^2 = 225 64 + 225 = 289